for 15 years, would like to buy his wife an expenisive diamond ring with a platinum setting on their 30-year wedding sary Assume that the cost of the ring will be $32,121 in 15 years. Kirk currently has $4,500 to invest. What annual rate of refurn must Kirk earn on his estment to aocumulate enouph money to pay for the ing? Cliak the table icon toview the PVI table The annual rate retu Krkmust earn n his investment is %. (Round tot e nearest integer)
Answer #1

Future value = Present value ( 1 + r)n

32,121 = 4,500(1 + r)15

7.138 = (1 + r)15

1.140001 = 1 + r

r = 0.14 or 14%

Annual rate will be 14%.

Keys to use in a financial calculator:

PV -4500

FV 32121

N 15

CPT I/Y